what volume is occupied by 19.6g of methane (ch4) and 27c and 1.94 atm
Answer #1

Solution:

Volume occupied is calculated from ideal gas equation,

PV = nRT

P = pressure = 1.94 atm

, V = volume

R =gas constant = 0.0821 atm L / mol

n = number of moles

T = temperature = 27 + 273 = 300 K

Therefore, V = nRT / P

Molar mass of methane = 16 g/ mol

Number of moles of methane = mass / molar mass

= 19.6 g / 16 g mol-1 = 1.225 mol

Hence, V = nRT /P

= 1.225 mol-1 x 0.0821 L atm mol-1 x 300 K / 1.94 atm

= 15.55 L

Hence, V = 15.55 L